If one leg of a triangle is 8 feet and one leg is 21 feet. What is the remaining leg?
Found 2 solutions by Mr.positive, JBarnum:
Answer by Mr.positive(13)

This exercise can't be answered as is. There is a mistake in how the problem is written! Are you sure that side that measures 21 feet is not the hypotenuse? Revise the exercise and I'll be happy to help you!
mathteacher2010@aol.com

Answer by JBarnum(2146)

um theres multiple answsers, technically 2 possible answers so you might want to reread the question to find out which side is the hypotnuse. if a= 8 and b=21 then C is greater than 21
if a=8 b=? c=21 so b would be less tha 21
8^2+21^2=c^2
64+441=c^2
505=c^2
22.4722=c
second one could be
8^2+b^2=21^2
64+b^2=441
b^2=377
b=19.4165